OpenStudy (anonymous):

A bacteria culture begins with 13 bacteria that triple in amount at the end of every hour. How many bacteria exist in the culture at the end of the 10th hour? a. 13,312 b. 26,624 c. 767,637 d. 2,302,911

OpenStudy (anonymous):

c

OpenStudy (anonymous):

yea but how, i tried and got 255879

OpenStudy (anonymous):

#of bacteria = 13(3)^n where 'n' is the number of years. so 13(3)^10 = 767637

OpenStudy (anonymous):

no i thought it was 13(3)^(10-1)

OpenStudy (anonymous):

Why would you do that?

OpenStudy (anonymous):

being that the formula is an = a1(r)^(n-1)

OpenStudy (anonymous):

That's for geometric sequence. This is an exponential equation, not a geometric sequence.

OpenStudy (anonymous):

oh so how do u no what is different

OpenStudy (anonymous):

General formula--> c= a(b)^n It's like the compound interest formula. a = initial amount/money. b = rate of increase/how many times it get's multiplied. n = number of times it happens/number of years.

OpenStudy (anonymous):

oh okay, i hope ur right... thank u!

OpenStudy (anonymous):

Anytime.

OpenStudy (anonymous):

Was correct, ty!